N = 774958P96Q If P= 2 and the number N is divisible by 3 ,then numbef of possible values of Q is/ are (1) 0 (2) 2 (3)3 (4)4

N = 774958P96Q 
If P= 2 and the number  N is divisible by 3 ,then numbef of possible values of Q is/ are
(1) 0  
(2) 2
(3)3
(4)4

Grade:11

2 Answers

Nishant Vora IIT Patna
7 years ago
774958P96Q

Sum of digits = 55 + P + Q = 57 +Q

So Q = 0,3,6,9
Ujjwal
3 years ago
If p=2
N=774958296Q
Divisibility rule of 3
Sum of digits of number should be divisible by 3
57+Q should be divisible by 3
Possible values=57,60,63,66
Q=0,3,6,9